# Explore Spurious Correlations at the Concept Level in Language Models for Text Classification

### 06 Jan 2024 v2

@@ 4.2 Spurious Correlations in ICL
After we find that LMs utilize the concept shortcuts in the fine-tuning data, we also want to ask whether LMs use the shortcuts in the exemplars of the prompts when performing ICL. As discussed in Section 2.4, for the concept $c$ in four datasets, we construct a prompt with 8 exemplars. For the concept-biased prompt $P_{biased}$ in all concepts except for “size” in Amazon Shoe and “service” in Yelp, 4 of the exemplars with concept $c$ are positive and 4 of the exemplars without concept $c$ are negative, and for the other two concepts, the labels are flipped. For the balanced prompt $P_{balanced}$, the label is evenly distributed for the exemplars with or without the concept. With the bias in label distribution for both texts with and without the concept, we expect the LM to use two types of shortcuts: 1. from texts with the concept $c$ to one sentiment and 2. from texts without the concept to the other sentiment. We evaluate the bias and utility performance for the ICL and present results in Table 4.
@@
For utility performance, when changing from $P_{balanced}$ to $P_{biased}$, the average Acc@C and Acc@NoC decrease from 76.80% to 76.42%, and from 76.37% to 75.58%, respectively, which means that spurious correlations harm utility performance regardless of the presence of concepts. For both Bias@C and accuracy, the relative change of two metrics in the ICL scenario is less than the fine-tuning setting. We conjecture that only a few exemplars in the prompts make it hard to form a strong shortcut inside the LMs between conceptual contents and a specific label, due to the pre-training of LMs on a large corpus.
@@ 5.1 Mitigation via Rebalancing
We consider two lines of existing work to mitigate spurious correlations: remove spurious components and rebalance the training dataset McCoy et al. (2019); Wang et al. (2022). Since it is challenging to identify the conceptual contents in each sentence to eliminate the shortcuts, we apply the dataset rebalancing to mitigate spurious correlations at the concept level.
The first method is to perform downsampling to create a dataset $\mathcal{D}^{c}_{down-bal}$ with a balanced label distribution under the concept $c$. The shortcoming of the method is that, for a highly biased dataset, it filters out a large proportion of examples with the majority labels, leading to a sacrifice of the utility performance. To solve the challenge, we propose an upsampling method using ChatGPT to generate counterfactual examples with minority labels.
Suppose that we need $\{a_{0},\cdots,a_{n}\}$ number of examples for labels $\{0,\cdots,n\}$ to make a balanced subset for texts with concept $c$ ($a_{i}\geq 0$). We first sample $a_{0}$ to $a_{n}$ numbers of examples from texts with labels 00 to $n$ but without concept $c$. Then we ask ChatGPT, $M_{a}$, to inject the concept $c$ into the selected texts while maintaining the sentiment. Given the input text $x^{\prime}$ without concept $c$, the injection prompt $P_{i}$ with the instruction, and the exemplars with concept $c$, the concept injection process is $x_{c}=M_{a}(P_{i}\|c\|x^{\prime})$, where $x_{c}$ the generated counterfactual input for concept $c$ and input $x^{\prime}$. We iteratively generate the counterfactual input $x_{c}$ and add it into the dataset $\mathcal{D}_{ori}$ to form a balanced dataset $\mathcal{D}^{c}_{up-bal}$ for concept $c$ with the upsampling method.
We also propose a baseline method that masks out words highly associated with the concept. This method is also used to verify whether the observed concept bias is equivalent to the word bias. We select words with the top 10 PMI for each concept (word examples for the Amazon shoe are in Table 2) and replace the selected words to the [MASK] token. We name the masked dataset as $\mathcal{D}^{c}_{mask}$.
@@ 5.2 Results of Mitigation Methods
To evaluate the effectiveness of proposed methods, we first select the concept with Bias@C > 1 in Table 3 for experiments and then perform the fine-tuning on three de-biased datasets $\mathcal{D}^{c}_{down-bal}$, $\mathcal{D}^{c}_{up-bal}$, and $\mathcal{D}^{c}_{mask}$. We report Bias@C and accuracy for DistilBERT in Table 5.
Upsampling method reduces the bias and increases utility performance  From Table 5, we observe that both data rebalancing methods are effective in mitigating spurious correlations in fine-tuned models. For the downsampling method ($\mathcal{D}^{c}_{down-bal}$), it mitigates the mean absolute values of Bias@C from 4.90% to 3.43%, compared to trained on $\mathcal{D}_{ori}$. However, for utility performance, the downsampling method obtains less accuracy in 4/8 cases for Acc@C and 5/8 cases for Acc@NoC, indicating that loss of data sacrifices utility. For the proposed upsampling method ($\mathcal{D}^{c}_{up-bal}$), the mean absolute values of Bias@C are effectively reduced from 4.90% to 2.74%. Furthermore, the average robust accuracy of Acc@C increases from 79.24% to 80.38% and the values of the average Acc@NoC are comparable (from 78.90% to 79.00%). This observation suggests that adding counterfactual concept texts to rebalance the data can reduce spurious correlations in the concept level, and more data involved in the fine-tuning can boost the models’ utility performance.
@@
In summary, among the three mitigation methods, adding the LLM-generated counterfactual inputs into the training data achieves the best performance in both the bias mitigation and utility aspects. We repeat three mitigation methods on the concept with Bias@C > 1 for the LLAMA2 models and present the results in Table 7 in Appendix B. We can observe similar patterns for the LLAMA2, which verifies the generalizability of our methods.
@@ 6 Related Work
Current work on studying spurious correlations and shortcut learning for pre-trained models can be split into two categories: utilize the shortcuts during training or during ICL (without parameter updates). For shortcut learning in the training dataset, a series of works explores how models take shortcuts in the data Tu et al. (2020); Sagawa et al. (2020); Geirhos et al. (2020); Ribeiro et al. (2020); Kaushik et al. (2019) and which aspects of shortcuts will be taken for the predictions in different NLP tasks McCoy et al. (2019); Jia and Liang (2017); Lai et al. (2021); Zhao et al. (2018); Niu et al. (2020); Poliak et al. (2018), leading to low generalization in the out-of-distribution data or in the designed adversarial data.
Due to the increasing development of LLM on ICL, researchers find that the design of the prompt significantly decides the LLM predictions Brown et al. (2020); Gao et al. (2020); Liu et al. (2023a); Zhou et al. (2023); Schick and Schütze (2020). Another line of work finds that LLMs are sensitive to a certain aspect of prompts and not robust when injecting adversarial triggers into prompt Lu et al. (2021); Zhao et al. (2021); Tang et al. (2023); Si et al. (2023); Zheng et al. (2023). Tang et al. (2023) shows that LLMs use multiple types of shortcuts in the prompts, from letters to words to text style, and Si et al. (2023) find that LLMs exhibit clear feature biases under the unspecified prompts. Our work focuses on assessing whether the models use the shortcuts in the general concept level, not covered by the previous work in fine-tuning or ICL aspects.
@@ 7 Conclusions
In this paper, we explore the spurious correlation in LMs at the general concept level in both fine-tuning and ICL settings. We find that LMs utilize the concept shortcut during training (or in demonstrations) when making predictions on unseen data, and more biased training data (or prompts) lead to more biased predictions. To mitigate the learned shortcut, we propose a rebalancing method by adding counterfactual examples generated by ChatGPT to the training data, which is shown to be effective through extensive experiments. Our work indicates that, besides the sensitive demographic concepts, LMs also form strong spurious correlations on the general concepts, encouraging future work to pay attention to unintended shortcut learning.
@@ 8 Limitations
Our work focuses on four classification datasets. Two of them are binary classification tasks, and two are multiclass classification tasks. We apply the difference in accuracy for different groups to measure the bias at the concept level. Future work could extend our framework and generalize it to more complex tasks, such as text generation.
@@ Appendix B Prompt Details and Supplementary Results
In Table 6 and Table 7, we repeat the experiments for fine-tuning LLAMA2 7B models for Section 4.1 and 5.2. In Table 8, 9, and 10, we present the details of the prompts in the annotation of the concept, the ICL experiments, and the countergactual sentence generation.

### 21 Feb 2024 v3

@@ 4.2 Spurious Correlations in ICL
As LMs exhibit clear evidence of utilizing the concept shortcuts in the fine-tuning data, we also want to ask whether LMs use the shortcuts in the exemplars of the prompts when performing ICL. As discussed in Section 2.4, for each concept $c$ in five datasets, we construct a prompt with eight exemplars. Following a similar setup for fine-tuning, we only include the majority class (positive or negative) for exemplars with concept $c$. Specifically, for “size” in Amazon Shoe and “service” in Yelp, four exemplars with concept $c$ have negative labels and the other four without concept $c$ have positive labels. The labels are flipped for the rest of the concepts. For the balanced prompt $P_{balanced}$, the label is evenly distributed for the exemplars with or without the concept. With the bias in label distribution for both texts with and without the concept, we expect the LM to use two types of shortcuts: a) from texts with the concept $c$ to one sentiment and b) from texts without the concept to the other sentiment. We present the utility and bias results of ICL for sentiment classification dataset in Table 4 and BoolQ in Table 6 (Appendix B).
@@
For utility performance, when changing from $P_{balanced}$ to $P_{biased}$, the average Acc@C and Acc@NoC decrease from 76.80% to 76.42% and from 76.37% to 75.58%, respectively, which means that spurious correlations harm utility performance regardless of the presence of concepts. For both Bias@C and accuracy, the relative change in the ICL scenario is less than the fine-tuning setting. We conjecture that a few exemplars in prompts make it hard to form a strong shortcut inside the LMs between conceptual contents and a specific label.
@@ 5.1 Mitigation via Rebalancing
We consider two lines of existing data-centric work to mitigate spurious correlations: remove spurious components and rebalance the training dataset McCoy et al. (2019); Wang et al. (2022). Since it is challenging to identify the conceptual contents in each sentence, we apply dataset rebalancing methods to mitigate the bias at the concept level.
We first downsample the dataset to achieve a balanced label distribution with respect to concept $c$, denoted as $\mathcal{D}^{c}_{down-bal}$. The shortcoming of the method is that, for a highly biased dataset, it filters out a large proportion of examples with the majority labels, leading to a sacrifice of the utility performance. To address this, we propose an upsampling method using ChatGPT to generate counterfactual examples with minority labels. The resulting dataset is denoted as $\mathcal{D}^{c}_{up-bal}$.
Suppose that we need $\{a_{0},\cdots,a_{n}\}$ number of examples for labels $\{0,\cdots,n\}$ to make a balanced subset for texts with concept $c$. We first sample $a_{0}$ to $a_{n}$ numbers of examples from texts with labels 00 to $n$ but without concept $c$. Then we ask ChatGPT, $M_{a}$, to inject the concept $c$ into the selected texts while maintaining the sentiment or the answer to questions. Given the input text $x^{\prime}$ without concept $c$, the injection prompt $P_{i}$ with the instruction, and the exemplars $h_{c}$ with concept $c$, the concept injection process is $x_{c}=M_{a}(P_{i}\|h_{c}\|x^{\prime})$, where $x_{c}$ is the generated counterfactual for concept $c$ and input $x^{\prime}$. We iteratively generate the counterfactual input $x_{c}$ and add it into the dataset $\mathcal{D}_{ori}$ to form a balanced dataset $\mathcal{D}^{c}_{up-bal}$ with upsampling.
We also propose a baseline method that masks out words highly associated with the concept. This method is used to verify whether balancing distributions of a few tokens removes conceptual shortcuts. We replace words with the top 10 PMI for each concept (word examples are in Table 2) to the [MASK] token and name the masked dataset as $\mathcal{D}^{c}_{mask}$.
@@ 5.2 Results of Mitigation Methods
To evaluate the effectiveness of proposed methods, we select concepts with Bias@C greater than 1 in Table 3 and fine-tune on three de-biased datasets $\mathcal{D}^{c}_{down-bal}$, $\mathcal{D}^{c}_{up-bal}$, and $\mathcal{D}^{c}_{mask}$. We report results for DistilBERT in Table 5 and 6 in Appendix B.
Upsampling method reduces the bias and increases utility performance  From Table 5, we observe that data rebalancing methods are effective in mitigating spurious correlations. For downsampling ($\mathcal{D}^{c}_{down-bal}$), it mitigates the mean absolute values of Bias@C from 4.90% to 3.43%, compared to trained on $\mathcal{D}_{ori}$. However, for utility performance, the downsampling obtains less accuracy in 4/8 cases for Acc@C and 5/8 cases for Acc@NoC, indicating that loss of data harms utility. For the upsampling method ($\mathcal{D}^{c}_{up-bal}$), the mean absolute values of Bias@C are effectively reduced from 4.90% to 2.74%. Furthermore, the average accuracy of Acc@C increases from 79. 24% to 80. 38%, and Acc@NoC is comparable. This observation suggests that adding counterfactual texts to rebalance the data can reduce spurious correlations in the concept level, and more data involved in the fine-tuning can boost the models’ utility performance.
@@
In summary, among the three mitigation methods, adding the LLM-generated counterfactual inputs achieves the best performance in both the bias mitigation and utility aspects. The same analysis on LLAMA2 models (Table 6 and 8 in Appendix B) reveals similar patterns, which shows the generalizability of our methods.
@@ 6 Related Work
Current work on studying spurious correlations for LMs can be split into two categories: utilize the shortcuts during training or ICL. For shortcut learning in training, a series of works explores how models take shortcuts in the data Tu et al. (2020); Sagawa et al. (2020); Geirhos et al. (2020); Ribeiro et al. (2020); Kaushik et al. (2019) and which aspects of shortcuts will be taken for the predictions in different NLP tasks McCoy et al. (2019); Jia and Liang (2017); Lai et al. (2021); Zhao et al. (2018); Niu et al. (2020); Poliak et al. (2018), leading to low generalization in the out-of-distribution data or in the designed adversarial data.
Due to the increasing development of LLM on ICL, researchers find that the design of the prompt significantly decides the LLM predictions Brown et al. (2020); Gao et al. (2020); Liu et al. (2023a); Zhou et al. (2023); Schick and Schütze (2020). Another line of work finds that LLMs are sensitive to a certain aspect of prompts and not robust when injecting adversarial triggers into prompt Lu et al. (2021); Zhao et al. (2021); Tang et al. (2023); Si et al. (2023); Zheng et al. (2023). Tang et al. (2023) shows that LLMs use multiple types of shortcuts in the prompts, from letters to words to text style, and Si et al. (2023) find that LLMs exhibit clear feature biases under the unspecified prompts. Previous work also develops multiple methods to identify the topic or concept of text input Li et al. (2024a); Abraham et al. (2022); Blei et al. (2003). However, our paper is the first to focus on assessing whether the models use shortcuts at the general concept level.
@@ 7 Conclusions
In this paper, we explore the spurious correlation at the general concept level in both fine-tuning and ICL settings. We find that LMs utilize the concept shortcut in training data (or in demonstrations) when inference on unseen data, and more biased training data (or prompts) lead to more biased predictions. To mitigate the learned shortcut, we propose a rebalancing method by adding counterfactual examples generated from ChatGPT to the training data, shown to be effective through extensive experiments. Our work indicates that LMs form strong spurious correlations on general concepts, encouraging future work to pay attention to unintended shortcut learning.
@@ 8 Limitations
Our work focuses on five classification tasks. Three of them are binary classification tasks, and two are multiclass classification tasks. We apply the difference in accuracy for different groups (positive and negative) to measure the bias at the concept level. Moreover, future work could extend our framework and generalize the measurement of concept bias to more complex tasks, such as the evaluation of LLM on QA tasks Li et al. (2024b).
@@ Appendix B Prompt Details and Supplementary Results
In Table 6, we perform the same analysis on the BoolQ question and answering dataset for all experiments (ICL and fine-tuning) in Section 4. In Table 7 and Table 8, we repeat the experiments for fine-tuning LLAMA2 7B models for Section 4.1 and 5.2. In Table 9, 10, and 11, we present the details of the prompts in the annotation of the concept, the ICL experiments, and the countergactual sentence generation.
Distilbert BoolQ Country (pos <neg) BoolQ History (pos >neg)
Bias@C Acc@NoC Acc@C Bias@C Acc@NoC Acc@C
Trained on $\mathcal{D}_{ori}$1.22 61.23 60.95 5.21 59.93 57.85
Trained on $\mathcal{D}^{c}_{biased}$18.90 55.92 55.46 50.63 57.37 55.22
Trained on $\mathcal{D}^{c}_{down-bal}$ 4.84 57.93 58.79 -8.95 59.87 58.60
Trained on $\mathcal{D}^{c}_{up-bal}$ 2.45 59.54 59.74 -0.85 60.13 59.70
Trained on $\mathcal{D}^{c}_{mask}$ 2.90 60.71 59.69 -1.55 60.94 60.84
LLAMA2 Bias@C Acc@NoC Acc@C Bias@C Acc@NoC Acc@C
Trained on $\mathcal{D}_{ori}$9.72 67.78 72.68 0.82 78.87 80.43
Trained on $\mathcal{D}^{c}_{biased}$9.67 67.23 70.86 2.73 75.86 78.79
Trained on $\mathcal{D}^{c}_{down-bal}$10.18 77.11 81.61 1.20 76.11 77.64
Trained on $\mathcal{D}^{c}_{up-bal}$7.81 77.30 78.16 -3.23 80.62 82.40
Trained on $\mathcal{D}^{c}_{mask}$10.55 77.53 79.10 -7.73 78.91 80.73
GPT3.5 ICL Bias@C Acc@NoC Acc@C Bias@C Acc@NoC Acc@C
ICL with $P_{balanced}$1.80 81.98 83.99 -3.88 83.12 83.90
ICL with $P_{biased}$2.90 82.93 83.11 -3.93 82.03 82.82
Table 6: Fine-tuning and ICL performance for all experiments in Section 4 on BoolQ dataset of DistilBert, LLAMA2 (fine-tuning) and GPT3.5 (ICL) models. The smaller absolute values of Bias@C (smaller bias) and larger values of Acc are in bold.